Shift report 20030221 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Feb 21 2003 - 00:02:14 EST

  • Next message: Apache: "Shift report 20030221 00:00-08:00"
    ------------------------------------------------------------
    Date: Fri Feb 21 00:02:09 2003, Shift: 16:00-24:00
    Supervisor: BHS, with: MM
    Log book#14 pages 28 - 28
    
    Shift Summary:
    ------------------------------------------------------------
    1: Machine down all shift, diverse problems. First with an RF  cavity, then with some blockage of yellow sector 2. This is still unresolved, but a fill is being attempted all the same.
    2: No runs apart from pedestals...
    3: D1 and D2 tripped around 19.00. D2 reset without complaining, but D1 needed Magnet Watch to get back up. 
    
    That's it for this shift...
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Fri Feb 21 2003 - 00:06:13 EST